OpenStudy (anonymous):

simplify by removing factors of 1: 187u^9x^8/45u^2x^6

OpenStudy (anonymous):

\[\frac{187u^9x^8}{45u^2x^6} =\frac{187\times u^9\times x^8}{45\times u^2\times x^6}\] \[=\frac{187\times u^{9-2}\times x^{8-6}}{45} =\frac{187\times u^{7}\times x^{2}}{45}\] \[=\frac{187 u^{7} x^{2}}{45} \] is the simplified form